OpenStudy (anonymous):

Will give medal(: Which of these square roots has an answer that is a real number: -41/2 or 4- 1/2? Why?

OpenStudy (anonymous):

Since \[\sqrt{\frac{-41}{2} } = i \sqrt{\frac{41}{2} } \] is an imaginary number so \[\sqrt { 4- \frac{1}{2}} =\sqrt { \frac{8-1}{2}} =\sqrt { \frac{7}{2}} \] is a real number